OpenStudy (anonymous):

Just for fun: Which is the best mathematical expression or number for you and why? For me e^(i*pi)

OpenStudy (anonymous):

8128

OpenStudy (anonymous):

fourier transforms.. cause i love to say it.. it sounds very sophisticated :D.. i like to pronounce it "few re air" transforms ( Like the girl maggie from transformers does)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

2=1

OpenStudy (anonymous):

There are so many good numbers you just can't chose one

OpenStudy (anonymous):

no ishaan, there are INFINITE good numbers :P

OpenStudy (anonymous):

I am just asking which is the number you like the most?

OpenStudy (anonymous):

I told you very difficult for me.

OpenStudy (anonymous):

We like the reverse -e^pi =1

OpenStudy (anonymous):

ipi

OpenStudy (anonymous):

216 = 6^3 = 125 + 64 +27 = 5^3 + 4^3 +3^3 Spooky!